Let x be a random variable that represents the Law School Aptitude Test (LSAT) score of a law school applicant. Then x has a normal distribution with mean μ=150 and standard deviation σ=10. Find the probability that the LSAT score of a randomly chosen law school applicant is at least 170. Hint: You must find P (x ≥ 170).
Answer #1

Solution :

Given ,

mean = = 150

standard deviation = = 10

P (x ≥ 170)= 1 - P(x<170 )

= 1 - P[(x -) / < (170-150) / 10]

= 1 - P(z <2 )

Using z table

= 1 -0.9772

= 0.0228

probability= 0.0228